# LLM Internal States Reveal Hallucination Risk Faced With a Query

### v1

@@ Abstract
The hallucination problem of Large Language Models (LLMs) significantly limits their reliability and trustworthiness. Humans have a self-awareness process that allows us to recognize what we don’t know when faced with queries. Inspired by this, our paper investigates whether LLMs can estimate their own hallucination risk before response generation. We analyze the internal mechanisms of LLMs broadly both in terms of training data sources and across 15 diverse Natural Language Generation (NLG) tasks, spanning over 700 datasets. Our empirical analysis reveals two key insights: (1) LLM internal states indicate whether they have seen the query in training data or not; and (2) LLM internal states show they are likely to hallucinate or not regarding the query. Our study explores particular neurons, activation layers, and tokens that play a crucial role in the LLM perception of uncertainty and hallucination risk. By a probing estimator, we leverage LLM self-assessment, achieving an average hallucination estimation accuracy of 84.32% at run time.
@@ 1 Introduction
Our work takes a step further by investigating whether LLM internal states have some indication of hallucination risk given queries and whether it can be reliably estimated even before the actual response generation (Fig. 1). We conduct a comprehensive analysis of the LLMs internal mechanisms both in terms of training data sources and across 15 diverse NLG tasks, spanning over 700 datasets through an in-depth analysis. We explore particular neurons, different activation layers, and tokens that play a crucial role in the LLM perception of uncertainty and hallucination risk. Employing a probing estimator Belinkov (2022) on the internal states associated with the queries, we validate their self-awareness and ability to indicate uncertainty in two aspects: (1) Whether they have seen the query in training data, achieving an accuracy of 80.28%. (2) Whether they are likely to hallucinate regarding the query, achieving an average estimation accuracy of 84.32% across 15 NLG tasks. We propose that understanding these representations could offer a proactive approach to estimating uncertainty, potentially serving as an early indicator for the necessity of retrieval augmentation Wang et al. (2023) or as an early warning system.
@@ 3.1 Problem Formulation
Suppose we have an LLM $f$ parameterized by $\theta$. It is able to gain internal states $I$ and generate response $r$ given user query $q$ represented as $I,r=f_{\theta}(q)$. We aim to investigate the self-awareness of LLM, specifically how their internal states $I$ relate to their level of hallucination risk $h$ when faced with a query $q$. This correlation is formulated by $u=\mathds{E}(I_{\theta,q})$, where $\mathds{E}$ signifies an estimator function. The combination of model-specific attribute via $f_{\theta}$, and query-only attribute via $q$, allows it to accurately capture the characteristics of individual LLMs and fosters a more efficient prediction mechanism that mirrors human cognitive processes.
We employ a dataset $\mathcal{D}_{\theta}=\{\langle I_{\theta,q,i}^{\text{train}},h^{\text{train}}_{i}\rangle\}_{i=1}^{N}$ consisting of $N$ query-label pairs. These pairs serve to represent the behavior of $f_{\theta}$ Here, $h^{\text{train}}_{i}$ denotes the hallucination risk label, which is determined based on (1) the query’s presence in the training data or (2) the degree of hallucination in the response $r$ to $q$. Thus, our objective is mathematically expressed as:
@@ 3.2 Data Construction
The unseen queries will trigger hallucinations due to the lack of information within the model’s training data when the model doesn’t refuse to respond. In other words, hallucinations triggered by unseen queries are data-related. To investigate the distinguishability between seen and unseen queries, we construct a compact dataset consisting of two distinct sets of queries. For the seen group, we collect historical BBC news in 2020 likely exposed during LLM’s training. For the unseen group, we collect recent BBC news in 2024 after the release of the LLM we analyze. To ensure comparability, we ensure these two sets share similar length distributions and semantic information via sentence embeddings  [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2.
@@
We first construct data using LLM to directly generate responses to queries in diverse NLG tasks. Subsequently, for labeling the responses and corresponding queries, a comprehensive integration of NLG metrics assesses the levels of hallucination. This generation only uses the parametric knowledge of LLM which is a proxy of performance in real-world applications. This generation process can also be in other settings, such as retrieval-augmented generation, to explore whether the internal states can estimate hallucination risk or other aspects’ performance in these settings.
@@ 3.3 Preliminary Analysis: Neurons for Hallucination Perception from Internal States
Internal states play a crucial role in language models, encapsulating rich contextual and semantic information learned from predicting tokens. They are adept at recognizing complex patterns and relationships pertinent to various NLP tasks, which positions them as potentially powerful tools for estimating the risk of hallucinations Azaria and Mitchell (2023); Liu et al. (2023); Chen et al. (2024). To analyze the self-assess sense of internal states and the role of specific neurons in the uncertainty and hallucination estimation, we employ a feature selection method based on Mutual Information Kraskov et al. (2004). This technique measures the relevance of different features for distinguishing between the categories in a dataset.
In the context of NLG tasks including dialogue, QA, and translation, we select the eight most significant neurons from the last activation layer and visualize them in Fig. 2. We observe that these neurons exhibit sensitivity to uncertainty, allowing them to distinguish between different hallucination levels given known and unknown queries. In other words, there exist individual neurons within LLM that can fairly perceive uncertainty and predict future hallucinations. This approach not only enhances our understanding of the neural correlates of hallucinations but also paves the way for developing targeted interventions that mitigate the effects of hallucinations.
@@ 3.4 Internal State-based Estimator
Based on the above preliminary analysis, we use internal states corresponding to the last token of queries, denoted as $x_{q}$. These representations, taken from a specified layer within the LLM, serve as the input to our estimation model. The accessibility and ease of obtaining these states further underscore their practicality for such applications.
@@ 4 Experiments
4.1 Dataset
(1) Seen/Unseen Query in Training Data
LatestEval Li et al. (2024) is a benchmark designed to tackle data contamination in evaluation through dynamic and time-sensitive construction. It includes data from various time periods. In our study, we utilize BBC news collected in Jan. 2020 and 2024, which correspond to periods before and after the release of the LLM. The selected dataset includes 2.8k samples allowing us to represent both seen and unseen data in our analysis.
(2) Hallucination Risk faced with the Query
Super-Natural Instructions Wang et al. (2022b) is a large and diverse benchmark including 1,616 diverse NLP datasets and their expert-written instructions. Among this benchmark, we select 15 NLG task categories including QA, Summarization, Translation, etc consisting of over 700 datasets. Please find the full list of NLG task categories in Fig. 3 or 4 and the full list of tasks in Tab. D in § A. We also leverage ANAH Ji et al. (2024) to test our estimator’s performance in the hallucination aspect and its generalization. ANAH is a bilingual dataset that offers analytical annotation of Hallucinations in LLMs within Generative QA. Our work uses English samples and treats the hallucination type as the label in the testing stage.
@@ 5.1 Results for Internal State-based Estimator
We evaluate our internal state-based estimator trained to distinguish unseen and seen questions. The F1 and accuracy scores reach 80.28% and 80.24% These high results shed light on the effectiveness of our internal state-based method in identifying unseen queries. This phenomenon is aligned with the previous works Kadavath et al. (2022); Yin et al. (2023b) which find the model can distinguish answerable and unanswerable questions that include future information.
@@ 5.2 Analysis
As shown in Tab. 1, we evaluate the generalization across different NLG tasks and within the same NLG task. Specifically, we examine zero-shot performance in QA and translation. While the zero-shot performance within these individual tasks is acceptable, the cross-task generalization remains relatively weak, aligned with the findings reported by Kadavath et al. (2022). In addition, we evaluate our estimator trained in QA on the out-of-domain hallucination QA dataset ANAH. The F1 score reaches 78.56% and the accuracy is 78.83%. These relatively high results shed light on the effectiveness of our internal state estimator in handling hallucination challenges and further show our generalization capabilities. Therefore, the features in internal states are shared with OOD data within the same task but not shared across tasks.
@@
More analysis is described in Appendix B including treating separate metrics as continuous regression labels and different estimator backbones.
@@ 6 Related Work
Researchers investigate the boundary of parametric knowledge in LLMs which aim to uncover what models know and where their capabilities end. Current research on the knowledge boundary predominantly narrows down this investigation to specialized tasks, charting a map of knowns and unknowns within these defined borders. In the representative QA field, Rajpurkar et al. (2018) and Yin et al. (2023b) collect datasets containing known and unknown questions and develop classifiers to differentiate them. The concept of unanswerability in these works is universal and model-agnostic, such as philosophical questions and unsolved mysteries. In contrast, Kadavath et al. (2022) focuses on model-specific prediction of which questions the model will be able to answer correctly. This approach requires direct fine-tuning of the LLM to make it a classifier in a few-shot setting. Out-of-domain or out-of-distribution detection Zhou et al. (2023); Ryu et al. (2018); Tan et al. (2019); Yang et al. (2021); Zheng et al. (2020) are also relevant areas dealing with the differentiation of unknown/unseen from training data, with main focus on classification tasks. Our method is versatile across various NLG tasks without requiring fine-tuning of LLMs.
@@
The phenomenon of hallucination in NLG encourages a variety of detection methods Min et al. (2023); Ji et al. (2024); Li et al. (2023); Scialom et al. (2021). Some of these methods delve into the internal states for detection. Azaria and Mitchell (2023), for example, collect a true-false statement dataset with artificial guidance and the classification results indicate that the LLMs’ internal state can reveal the truthfulness of statements. INSIDE Chen et al. (2024) also leverages LLMs’ internal states and proposes EigenScore for evaluating the self-consistency of responses, thereby serving as a proxy for hallucination levels. MIND Su et al. (2024), an unsupervised training approach, distinguishes the hallucinated continuation text from the original Wikipedia content based on internal states. On the other hand, Xiao and Wang (2021) shows evidence that higher uncertainty corresponds to a higher hallucination probability. Uncertainty estimation methods, such as  Xiao et al. (2022); Xiong et al. (2023); Kadavath et al. (2022), predict the reliability of their natural language outputs and can also serve as a tool for hallucination detection. Previous works leverage the LLM’s internal states for the text to be measured which is not necessary from the same LLM. Differently, this work focuses on self-awareness corresponding to the queries.
@@ 7 Conclusion
Inspired by human self-awareness, this work demonstrates the latent capacity of LLMs to self-assess and estimate hallucination risks prior to response generation. We conduct a comprehensive analysis of the internal states of LLMs both in terms of training data sources and across 15 NLG tasks with over 700 datasets. Employing a probing estimator on the internal states associated with the queries, we assess their self-awareness and ability to indicate uncertainty in two aspects: (1) recognizing whether they have seen the query in training data, achieving an accuracy of 80.28% [^3] Please refer to the first parts of § 3.2, § 4.1, and § 5.1.. (2) recognizing whether they are likely to hallucinate when faced with the query. The results demonstrate that internal state-based self-assessment outperforms PPL-based and prompting-based baselines, with an average estimation accuracy of 84.32% across all tested datasets. In addition, we explore the role of particular neurons in uncertainty and hallucination perception and reveal a positive correlation between the depth of activation layers in an LLM and its predictive accuracy. The consistency of internal states across different models suggests a potential for zero-shot transfer, but model-specific estimation is the optimal strategy. Challenges of generalizing these findings across different tasks are noted, despite observing promising generalizations within the same NLG tasks.
@@ Appendix D AI Assistants Using
In this paper, we use ChatGPT to improve the writing at grammar level.
@@
Data to Text 9 task1728 web nlg data to text, task1598 nyc long text generation, task1631 openpi answer generation, task677 ollie sentence answer generation, task957 e2e nlg text generation generate, task760 msr sqa long text generation, task1407 dart question generation, task102 commongen sentence generation, task1409 dart text generation
@@
Dialogue Generation 13 task574 air dialogue sentence generation, task361 spolin yesand prompt response classification, task576 curiosity dialogs answer generation, task1603 smcalflow sentence generation, task1714 convai3 sentence generation, task1730 personachat choose next, task565 circa answer generation, task611 mutual multi turn dialogue, task1729 personachat generate next, task1600 smcalflow sentence generation, task639 multi woz user utterance generation, task1590 diplomacy text generation, task360 spolin yesand response generation
@@
Explanation 6 task295 semeval 2020 task4 commonsense reasoning, task192 hotpotqa sentence generation, task593 sciq explanation generation, task1369 healthfact sentence generation, task223 quartz explanation generation, task134 winowhy reason generation
@@
Grammar Error Correction 2 task1415 youtube caption corrections grammar correction, task1557 jfleg answer generation
Number Conversion 2 task1703 ljspeech textmodification, task1704 ljspeech textmodification
@@
Overlap Extraction 2 task039 qasc find overlapping words, task281 points of correspondence
@@
Paraphrasing 12 task776 pawsx japanese text modification, task045 miscellaneous sentence paraphrasing, task770 pawsx english text modification, task771 pawsx korean text modification, task774 pawsx german text modification, task177 para-nmt paraphrasing, task466 parsinlu qqp text modification, task775 pawsx chinese text modification, task1614 sick text modify, task773 pawsx spanish text modification, task132 dais text modification, task772 pawsx french text modification
@@
Program Execution 90 task113 count frequency of letter, task1151 swap max min, task509 collate of all alphabetical and numerical elements in list separately, task100 concatenate all elements from index i to j, task096 conala list index subtraction, task365 synthetic remove vowels, task622 replace alphabets in a list by their position in english alphabet, task852 synthetic multiply odds, task1088 array of products, task1405 find median, task637 extract and sort unique digits in a list, task1446 farthest integers, task506 position of all alphabetical elements in list, task378 reverse words of given length, task093 conala normalize lists, task1404 date conversion, task097 conala remove duplicates, task372 synthetic palindrome numbers, task755 find longest substring and replace its sorted lowercase version in both lists, task636 extract and sort unique alphabets in a list, task267 concatenate and reverse all elements from index i to j, task162 count words starting with letter, task159 check frequency of words in sentence pair, task208 combinations of list, task1316 remove duplicates string, task504 count all alphabetical elements in list, task079 conala concat strings, task158 count frequency of words, task507 position of all numerical elements in list, task374 synthetic pos or neg calculation, task1087 two number sum, task163 count words ending with letter, task756 find longert substring and return all unique alphabets in it, task101 reverse and concatenate all elements from index i to j, task1551 every ith element from kth element, task606 sum of all numbers in list between positions i and j, task368 synthetic even or odd calculation, task1150 delete max min, task851 synthetic multiply evens, task377 remove words of given length, task063 first i elements, task064 all elements except first i, task245 check presence in set intersection, task161 count words containing letter, task605 find the longest common subsequence in two lists, task850 synthetic longest palindrome, task157 count vowels and consonants, task373 synthetic round tens place, task206 collatz conjecture, task1443 string to number, task123 conala sort dictionary, task244 count elements in set union, task499 extract and add all numbers from list, task124 conala pair averages, task1444 round power of two, task099 reverse elements between index i and j, task1089 check monotonic array, task1188 count max freq char, task125 conala pair differences, task488 extract all alphabetical elements from list in order, task1542 every ith element from starting, task1194 kth largest element, task371 synthetic product of list, task1406 kth smallest element, task095 conala max absolute value, task1315 find range array, task243 count elements in set intersection, task1331 reverse array, task062 bigbench repeat copy logic, task122 conala list index addition, task091 all elements from index i to j, task369 synthetic remove odds, task497 extract all numbers from list in order, task505 count all numerical elements in list, task205 remove even elements, task1189 check char in string, task1445 closest integers, task094 conala calculate mean, task160 replace letter in a sentence, task1148 maximum ascii value, task098 conala list intersection, task078 all elements except last i, task523 find if numbers or alphabets are more in list, task370 synthetic remove divisible by 3, task367 synthetic remove floats, task1190 add integer to list, task376 reverse order of words, task600 find the longest common substring in two strings, task207 max element lists, task366 synthetic return primes

### v2

@@ Abstract
The hallucination problem of Large Language Models (LLMs) significantly limits their reliability and trustworthiness. Humans have a self-awareness process that allows us to recognize what we don’t know when faced with queries. Inspired by this, our paper investigates whether LLMs can estimate their own hallucination risk before response generation. We analyze the internal mechanisms of LLMs broadly both in terms of training data sources and across 15 diverse Natural Language Generation (NLG) tasks, spanning over 700 datasets. Our empirical analysis reveals two key insights: (1) LLM internal states indicate whether they have seen the query in training data or not; and (2) LLM internal states show they are likely to hallucinate or not regarding the query. Our study explores particular neurons, activation layers, and tokens that play a crucial role in the LLM perception of uncertainty and hallucination risk. By a probing estimator, we leverage LLM self-assessment, achieving an average hallucination estimation accuracy of 84.32% at run time. [^1] The source code can be obtained from https://github.com/ziweiji/Internal_States_Reveal_Hallucination
@@ 1 Introduction
Our work takes a step further by investigating whether LLM internal states have some indication of hallucination risk given queries and whether it can be reliably estimated even before the actual response generation (Fig. 1). We conduct a comprehensive analysis of LLMs internal mechanisms in terms of training data sources and across 15 diverse NLG tasks that extend beyond the QA task Snyder et al. (2023); Slobodkin et al. (2023) and span over 700 datasets. We explore particular neurons, different activation layers, and tokens that play a crucial role in the LLM perception of uncertainty and hallucination risk. Employing a probing estimator Belinkov (2022) on the internal states associated with the queries, we validate their self-awareness and ability to indicate uncertainty in two aspects: (1) Whether they have seen the query in training data, achieving an accuracy of 80.28%. (2) Whether they are likely to hallucinate regarding the query, achieving an average estimation accuracy of 84.32% across 15 NLG tasks. We propose that understanding these representations could offer a proactive approach to estimating uncertainty, potentially serving as an early indicator for the necessity of retrieval augmentation Wang et al. (2023) or as an early warning system.
@@ 3.1 Problem Formulation
Suppose we have an LLM $f$ parameterized by $\theta$. It is able to gain internal states $I$ and generate response $r$ given user query $q$ represented as $I_{\theta,q},r_{\theta,q}=f_{\theta}(q)$. We aim to investigate the self-awareness of LLM, specifically how their internal states $I$ relate to their level of hallucination risk $h$ when faced with a query $q$.
We employ a dataset $\mathcal{D}_{\theta}=\{\langle I_{\theta,q,i}^{\text{train}},h^{\text{train}}_{i}\rangle\}_{i=1}^{N}$ consisting of $N$ query-label pairs. These pairs serve to represent the behavior of $f_{\theta}$ Here, $h^{\text{train}}_{i}$ denotes the level of hallucination risk, which is labeled based on (1) the query’s presence in the training data or (2) the degree of hallucination in the response $r$ to $q$. Thus, our objective is mathematically expressed as:
@@
Here, $\mathds{E}$ signifies an estimator function. The combination of model-specific attribute via $f_{\theta}$, and query-only attribute via $q$, allows it to accurately capture the characteristics of individual LLMs and fosters a more efficient prediction mechanism that mirrors human cognitive processes.
@@ 3.2 Data Construction
The unseen queries will trigger hallucinations due to the lack of information within the model’s training data when the model doesn’t refuse to respond. In other words, hallucinations triggered by unseen queries are data-related. To investigate the distinguishability between seen and unseen queries, we construct a compact dataset consisting of two distinct sets of queries. For the seen group, we utilize historical BBC news in 2020 highly likely exposed during LLM’s training. For the unseen group, we utilize recent BBC news in 2024 after the release of the LLM we analyze. To ensure comparability, we ensure these two sets share similar length distributions and semantic information via sentence embeddings  [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. The two groups are both from LatestEval Li et al. (2024), a benchmark designed to tackle data contamination in evaluation through dynamic and time-sensitive construction. The query is Tell more details about the news: {news_title}.
@@
We first construct data using LLM to directly generate responses to queries in diverse NLG tasks. Subsequently, for labeling the responses and corresponding queries, a comprehensive integration of NLG metrics assesses the levels of hallucination.
We select 15 NLG task categories including QA, Summarization, Translation, etc consisting of over 700 datasets from Super-Natural Instructions benchmark Wang et al. (2022b)  [^3] Please find the full list of NLG task categories in Fig. 3 or 4 and the full list of tasks in Tab. D in § A.. This generation only uses the parametric knowledge of LLM which is a proxy of performance in real-world applications. This generation process can also be in other settings, such as retrieval-augmented generation (RAG), to explore whether the internal states can estimate hallucination risk or other aspects’ performance in these settings.
@@ 3.3 Preliminary Analysis: Neurons for Hallucination Perception from Internal States
Internal states play a crucial role in language models, encapsulating rich contextual and semantic information learned from predicting tokens. They are adept at recognizing complex patterns and relationships pertinent to various NLP tasks, which positions them as potentially powerful tools for estimating the risk of hallucinations Azaria and Mitchell (2023); Liu et al. (2023); Chen et al. (2024). Furthermore, previous works Azaria and Mitchell (2023); Ahdritz et al. (2024); Liu et al. (2024) have demonstrated that the activations of the last token from the last layer in LLMs contain one of the most useful features. Therefore, we take these representations for preliminary analysis on the self-assess sense of internal states and the role of specific neurons in the uncertainty and hallucination estimation. Specifically, we employ a feature selection method based on Mutual Information Kraskov et al. (2004) to measure the relevance of different features/dimensions for distinguishing between the categories in a dataset.
In the context of NLG tasks including dialogue, QA, and translation, we select the eight most significant neurons/dimensions from the last activation layer and visualize them in Fig. 2. We observe that these neurons exhibit sensitivity to uncertainty, allowing them to distinguish between different hallucination levels given known and unknown queries. In other words, there exist individual neurons within LLM that can fairly perceive uncertainty and predict future hallucinations. This approach not only enhances our understanding of the neural correlates of hallucinations but also paves the way for developing targeted interventions that mitigate the effects of hallucinations.
@@ 3.4 Internal State-based Estimator
Based on the above preliminary analysis and previous works, we use the activations corresponding to the last token of queries from a specified layer in LLMs, denoted as $x_{q}$, as the input for our estimation model. The accessibility and ease of obtaining these states further underscore their practicality for such applications.
@@ 5.1 Results for Internal State-based Estimator
We evaluate our internal state-based estimator trained to distinguish unseen and seen questions. The F1 and accuracy scores reach 80.28% and 80.24%. These high results shed light on the effectiveness of our internal state-based method in identifying unseen queries. This phenomenon is aligned with the previous works Kadavath et al. (2022); Yin et al. (2023b) which find the model can distinguish answerable and unanswerable questions that include future information.
@@
More results are described in Appendix B including treating separate metrics (Rouge-L, NLI, and QuestEval) as continuous regression labels and different estimator backbones.
@@ 5.2 Analysis
As shown in Tab. 1, we evaluate the generalization across different NLG tasks and within the same NLG task. Specifically, we examine zero-shot performance in QA and translation. While the zero-shot performance within these individual tasks is acceptable, the cross-task generalization remains relatively weak, aligned with the findings reported by Kadavath et al. (2022).
In addition, we evaluate our estimator trained in QA on the out-of-domain hallucination QA dataset ANAH Ji et al. (2024) to test our estimator’s performance in the hallucination aspect and its generalization. ANAH is a bilingual dataset that offers analytical annotation of Hallucinations in LLMs within Generative QA. Our work uses English samples and treats the hallucination type as the label in the testing stage. The F1 score reaches 78.56% and the accuracy is 78.83%. These relatively high results shed light on the effectiveness of our internal state estimator in handling hallucination challenges and further show our generalization capabilities. Therefore, the features in internal states are shared with OOD data within the same task but not shared across tasks.
@@ 6 Related Work
Researchers investigate the boundary of parametric knowledge in LLMs which aim to uncover what models know and where their capabilities end. Current research on the knowledge boundary predominantly narrows down this investigation to specialized tasks, charting a map of knowns and unknowns within these defined borders. In the representative QA field, Rajpurkar et al. (2018) and Yin et al. (2023b) collect datasets containing known and unknown questions and develop classifiers to differentiate them. The concept of unanswerability in these works is universal and model-agnostic, such as philosophical questions and unsolved mysteries. In contrast, Kadavath et al. (2022); Slobodkin et al. (2023), and Gottesman and Geva (2024) focus on model-specific prediction for questions. The former focuses on whether the model will answer correctly, while the latter focuses on whether the model linguistically refuses to answer. Out-of-domain or out-of-distribution detection Zhou et al. (2023); Ryu et al. (2018); Tan et al. (2019); Yang et al. (2021); Zheng et al. (2020) are also relevant areas dealing with the differentiation of unknown/unseen from training data, with main focus on classification tasks. Our method is versatile across various NLG tasks without requiring fine-tuning of LLMs.
@@
The phenomenon of hallucination in NLG encourages a variety of detection methods Min et al. (2023); Ji et al. (2024); Li et al. (2023); Scialom et al. (2021). Some of these methods delve into the internal states for detection. Azaria and Mitchell (2023), for example, collect a true-false statement dataset with artificial guidance and the classification results indicate that the LLMs’ internal state can reveal the truthfulness of statements. INSIDE Chen et al. (2024) also leverages LLMs’ internal states and proposes EigenScore for evaluating the self-consistency of responses, thereby serving as a proxy for hallucination levels. MIND Su et al. (2024), an unsupervised training approach, distinguishes the hallucinated continuation text from the original Wikipedia content based on internal states. Snyder et al. (2023) explore the query-only detection within the QA task based on internal states, gradients, and probabilities. On the other hand, Xiao and Wang (2021) shows evidence that higher uncertainty corresponds to a higher hallucination probability. Uncertainty estimation methods, such as  Xiao et al. (2022); Xiong et al. (2023); Kadavath et al. (2022), predict the reliability of their natural language outputs and can also serve as a tool for hallucination detection. Previous works leverage the LLM’s internal states for the text to be measured which is not necessary from the same LLM. Differently, this work focuses on self-awareness corresponding to the queries across multiple NLG tasks.
@@ 7 Conclusion
Inspired by human self-awareness, this work demonstrates the latent capacity of LLMs to self-assess and estimate hallucination risks prior to response generation. We conduct a comprehensive analysis of the internal states of LLMs both in terms of training data sources and across 15 NLG tasks with over 700 datasets. Employing a probing estimator on the internal states associated with the queries, we assess their self-awareness and ability to indicate uncertainty in two aspects: (1) recognizing whether they have seen the query in training data, achieving an accuracy of 80.28% [^5] Please refer to the first parts of § 3.2 and § 5.1.. (2) recognizing whether they are likely to hallucinate when faced with the query. The results demonstrate that internal state-based self-assessment outperforms PPL-based and prompt-based baselines, with an average estimation accuracy of 84.32% across all tested datasets. In addition, we explore the role of particular neurons in uncertainty and hallucination perception and reveal a positive correlation between the depth of activation layers in an LLM and its predictive accuracy. The consistency of internal states across different models suggests a potential for zero-shot transfer, but model-specific estimation is the optimal strategy. Challenges of generalizing these findings across different tasks are noted, despite observing promising generalizations within the same NLG tasks.
@@ 9 Ethical Considerations
Acknowledgement
This work has been supported by the China NSFC Project NSFC21EG14 (No. 62120106006), SAAIR Project (No. Z1286), HKJCCT21EG01 (RG192), and Care-E Project (FS116).
@@ Appendix D AI Assistants Using
In this paper, we use ChatGPT to improve the writing at the grammar level.
@@
Data to Text 9 Task 1728: web nlg data to text, Task 1598: nyc long text generation, Task 1631: openpi answer generation, Task 677: ollie sentence answer generation, Task 957: e2e nlg text generation generate, Task 760: msr sqa long text generation, Task 1407: dart question generation, Task 102: commongen sentence generation, Task 1409: dart text generation
@@
Dialogue Generation 13 Task 574: air dialogue sentence generation, Task 361: spolin yesand prompt response classification, Task 576: curiosity dialogs answer generation, Task 1603: smcalflow sentence generation, Task 1714: convai3 sentence generation, Task 1730: personachat choose next, Task 565: circa answer generation, Task 611: mutual multi turn dialogue, Task 1729: personachat generate next, Task 1600: smcalflow sentence generation, Task 639: multi woz user utterance generation, Task 1590: diplomacy text generation, Task 360: spolin yesand response generation
@@
Explanation 6 Task 295: semeval 2020 Task 4: commonsense reasoning, Task 192: hotpotqa sentence generation, Task 593: sciq explanation generation, Task 1369: healthfact sentence generation, Task 223: quartz explanation generation, Task 134: winowhy reason generation
@@
Grammar Error Correction 2 Task 1415: youtube caption corrections grammar correction, Task 1557: jfleg answer generation
Number Conversion 2 Task 1703: ljspeech textmodification, Task 1704: ljspeech textmodification
@@
Overlap Extraction 2 Task 039: qasc find overlapping words, Task 281: points of correspondence
@@
Paraphrasing 12 Task 776: pawsx japanese text modification, Task 045: miscellaneous sentence paraphrasing, Task 770: pawsx english text modification, Task 771: pawsx korean text modification, Task 774: pawsx german text modification, Task 177: para-nmt paraphrasing, Task 466: parsinlu qqp text modification, Task 775: pawsx chinese text modification, Task 1614: sick text modify, Task 773: pawsx spanish text modification, Task 132: dais text modification, Task 772: pawsx french text modification
@@
Program Execution 90 Task 113: count frequency of letter, Task 1151: swap max min, Task 509: collate of all alphabetical and numerical elements in list separately, Task 100: concatenate all elements from index i to j, Task 096: conala list index subtraction, Task 365: synthetic remove vowels, Task 622: replace alphabets in a list by their position in english alphabet, Task 852: synthetic multiply odds, Task 1088: array of products, Task 1405: find median, Task 637: extract and sort unique digits in a list, Task 1446: farthest integers, Task 506: position of all alphabetical elements in list, Task 378: reverse words of given length, Task 093: conala normalize lists, Task 1404: date conversion, Task 097: conala remove duplicates, Task 372: synthetic palindrome numbers, Task 755: find longest substring and replace its sorted lowercase version in both lists, Task 636: extract and sort unique alphabets in a list, Task 267: concatenate and reverse all elements from index i to j, Task 162: count words starting with letter, Task 159: check frequency of words in sentence pair, Task 208: combinations of list, Task 1316: remove duplicates string, Task 504: count all alphabetical elements in list, Task 079: conala concat strings, Task 158: count frequency of words, Task 507: position of all numerical elements in list, Task 374: synthetic pos or neg calculation, Task 1087: two number sum, Task 163: count words ending with letter, Task 756: find longert substring and return all unique alphabets in it, Task 101: reverse and concatenate all elements from index i to j, Task 1551: every ith element from kth element, Task 606: sum of all numbers in list between positions i and j, Task 368: synthetic even or odd calculation, Task 1150: delete max min, Task 851: synthetic multiply evens, Task 377: remove words of given length, Task 063: first i elements, Task 064: all elements except first i, Task 245: check presence in set intersection, Task 161: count words containing letter, Task 605: find the longest common subsequence in two lists, Task 850: synthetic longest palindrome, Task 157: count vowels and consonants, Task 373: synthetic round tens place, Task 206: collatz conjecture, Task 1443: string to number, Task 123: conala sort dictionary, Task 244: count elements in set union, Task 499: extract and add all numbers from list, Task 124: conala pair averages, Task 1444: round power of two, Task 099: reverse elements between index i and j, Task 1089: check monotonic array, Task 1188: count max freq char, Task 125: conala pair differences, Task 488: extract all alphabetical elements from list in order, Task 1542: every ith element from starting, Task 1194: kth largest element, Task 371: synthetic product of list, Task 1406: kth smallest element, Task 095: conala max absolute value, Task 1315: find range array, Task 243: count elements in set intersection, Task 1331: reverse array, Task 062: bigbench repeat copy logic, Task 122: conala list index addition, Task 091: all elements from index i to j, Task 369: synthetic remove odds, Task 497: extract all numbers from list in order, Task 505: count all numerical elements in list, Task 205: remove even elements, Task 1189: check char in string, Task 1445: closest integers, Task 094: conala calculate mean, Task 160: replace letter in a sentence, Task 1148: maximum ascii value, Task 098: conala list intersection, Task 078: all elements except last i, Task 523: find if numbers or alphabets are more in list, Task 370: synthetic remove divisible by 3, Task 367: synthetic remove floats, Task 1190: add integer to list, Task 376: reverse order of words, Task 600: find the longest common substring in two strings, Task 207: max element lists, Task 366: synthetic return primes